log(16)x=-1/2...how to solve for x
Answer by swincher4391(726)

The exponentiate (by base 16) both sides of the equation:
16^(log(16)(x)) = 16^(-1/2)
x = 16^(-1/2)
x = 1/(16^(1/2))
x = 1/sqrt(16)
x = 1/4 <-------Answer